Shift report 20030306 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Mar 06 2003 - 16:14:24 EST

  • Next message: Flemming Videbaek: "RUN update"
    ------------------------------------------------------------
    Date: Thu Mar  6 16:14:23 2003, Shift: 08:00-16:00
    Supervisor: ykl, with: rl
    Log book#14 pages 48 - 49
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam until 11:00. One hour of IR access. No beam until 
    the end of shift.
    
    2.
    MRS @90, A1050; FS @4, 1/1B    
    8219 Trig 3 42K, trig 2, 11K. 
    8220 Trig 3 30K, trig 2,  9K. 
    8221 Trig 3 18K, trig 2,  2K  
    8222 nonTPC pedestal, 8223 TPC pedestal
    
    3.During the access HI light sealed tygon supply tube to C4;
    and DB reconnected TOFII and replaced a discriminator.
    
    Plan: MRS @90 deg A1050 for "lots of statistics".
          FS @3deg 1/4B for 300K.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu Mar 06 2003 - 16:19:14 EST